OpenStudy (anonymous):

Write the equation y - 2 = 2(x - 3) in standard form. y - 2x = 4 2x - y = -6 2x + y = 6 2x - y = 4 Write the equation y - 2 = 2(x - 3) in standard form. y - 2x = 4 2x - y = -6 2x + y = 6 2x - y = 4 @Mathematics

OpenStudy (anonymous):

2x - y = 4

OpenStudy (anonymous):

2x-y=4

jhonyy9 (jhonyy9):

so y=2x-1 y=2x+4 y=2x+6 y= -2x+6 y=2x-4